solve the linear system:
2x = 3y
x+3y = 9

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

solve the linear system:
2x = 3y
x+3y = 9
--------------
Modify:
x = (3/2)y
---
Substitute for "x" and solve for "Y":
(3/2)y + 3y = 9
(9/2)y = 9
(1/2)y = 1
y = 2
-----
Since x = (3/2)y, x = (3/2)2 = 3
------------------------------------
Solution: x=3;y=2
======================
Cheers,
Stan H.

Answer by richwmiller(17219)

x+3y=9
3y=9-x
2x=3y
2x=9-x
3x=9
x=3
3+3y=9
3y=6
y=2